A thin glass prism of refractive index 1.5 is immersed in water of refractive index 1.3. If the angle of deviation in air for a particular ray be D,then that in water is

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
 
Let the angle of prism be A 
Deviation =(u-1)A
4°=(1.5-1)A
So we get A=8° now when we dip in water so refractive index of water w.r.t glass ug/uw=3/2/4/3=9/8
Deviation =(9/8-1)8=1°
Hence answer=1°
 
Regards
Arun